# Continual Release Moment Estimation with Differential Privacy

### v1

@@ 3.3 Comparison with Alternative Techniques
Besides JME, a number of alternatives methods are possible that could be used to privately estimate the first and second moments. In this section, we introduce some of them and discuss their relation to JME.
3.3.1 Independent Moment Estimation (IME)
A straight-forward way to privately estimate first and second moments is to estimate and privatize both of them separately, where the necessary amount of noise is determined by the composition theorem of the Gaussian mechanism (Abadi et al., 2016) (see Algorithm 4 in the appendix).
@@ 3.3.2 Concatenate-and-Split (CS)
When applicable, CS is easy to implement and produces unbiased moment estimates. However, like IME, it has an unfavorable privacy-accuracy trade-off curve compared to JME, as formalized in the following theorem.
Figure 1: Approximation errors for the first and second moments ($d=10$, $n=100$, $C_{1}=C_{2}=I,A=E_{1}$, $\zeta=1$, and $\sigma=1$) across three methods: $\lambda$-JME with $\lambda$, IME with varying $\alpha$ parameters, and CS with varying $\tau$ parameter. One can see that $\lambda$-JME Pareto-dominates the other two methods.
@@
The proof can be found in the appendix. Figure 1 visualizes these cases (Theorems 3.6 and 3.7) graphically.
3.3.3 Post-processing (PP)
Post-processing (PP) is another easy-to-use method for joint moment estimation. It has appeared in the literature (Sheffet, 2019), at least in its naive form without the matrix factorization mechanism. For any $x_{i}$, it first computes a private estimate $\widehat{x_{t}}$ by adding sufficient noise to it. It then sets $\widehat{x_{t}\otimes x_{t}}:=\widehat{x_{t}}\otimes\widehat{x_{t}}$, which is automatically private by the postprocessing property of DP, and uses it to estimate the second moment matrix without additional privacy protection. Algorithm 6 in the appendix provides pseudocode.
(a) $d=1$, $C_{1}=C_{2}=\text{I}$
(b) $d=1$, $C_{1}=C_{2}=A^{1/2}$
(c) $d=1000$, $C_{1}=C_{2}=\text{I}$
(d) $d=1000$, $C_{1}=C_{2}=A^{1/2}$
@@ 3.3.3 Post-processing (PP)
For given privacy parameters, PP and JME use the same noise strength to privatize the first moment and therefore their estimates are of equal quality. For the second moment, PP and JME differ in their characteristics between the low privacy (small noise variance) and the high privacy (large noise variance) regime. To allow for a quantitative comparison, we derive characterizations of the approximation quality of PP.
@@
Assume the same setting as for Theorem 3.3, except that we compute the estimates $\widehat{Y}$ and $\widehat{S}_{\text{PP}}$ with the PP method. Then, the expected approximation error of the second moment satisfies:
$\sup_{X\in\mathcal{X}}\mathbb{E}\|S-\widehat{S}_{\text{PP}}\|^{2}_{\text{F}}=\overbrace{d\sigma^{4}\zeta^{2}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})}^{\text{bias}}$
$\ \ +d(d+1)\sigma^{4}\zeta^{2}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))$ (8)
$\ \ +2(d+1)\sigma^{2}\zeta^{2}\|C_{1}\|_{1\to 2}^{2}\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{\top}A_{2}\circ Q)XX^{T}),$
where $Q=C_{1}^{-1}C_{1}^{-\top}$ and $E_{Q}=\operatorname{\operatorname{diag}}(Q)\operatorname{\operatorname{diag}}(Q)^{\top}$. For debiased PP, the term marked ”bias” does not occur.
@@
Assume that JME and PP use trivial factorizations, i.e. $C_{1}=C_{2}=I$. Then, JME’s expected approximation error for the second moment is
$\mathbb{E}\|S-\widehat{S}\|^{2}_{\text{F}}$ $=4c_{d}d^{2}\sigma^{2}\zeta^{2}\|A_{2}\|_{\text{F}}^{2}$ (9)
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}\|S-\widehat{S}_{\text{PP}}\|^{2}_{\text{F}}$ $=\big{(}d(d+1)\sigma^{4}+2(d+1)\sigma^{2}\big{)}\zeta^{2}\|A_{2}\|_{\text{F}}^{2}$
$\quad+\underbrace{d\sigma^{4}\zeta^{2}\|\sum_{i}(A_{2})_{[i,\cdot]}\|_{2}^{2}}_{\text{bias}},$ (10)
where $\sum_{i}(A_{2})_{[i,\cdot]}$ is the row-wise summation of $A_{2}$.
For debiased PP, the term marked ”bias” does not occur.
Proof. The proofs follow directly from Theorem 3.3 and Lemma 3.8 by observing that $\|C_{1}\|_{1\to 2}=\|C_{2}\|_{1\to 2}=1$.
In dimension $d=1$, JME has an error of $c_{1}\sigma^{2}\|A_{2}\|_{\text{F}}^{2}$ versus $(2\sigma^{4}+4\sigma^{2})\|A_{2}\|_{\text{F}}^{2}$ for (debiased) PP. Since $c_{1}<1$, for one-dimensional data, JME’s error is always lower than PP’s (see Figure 2 for visual confirmation). In high dimensions ($d\gg 1$), the terms quadratic in $d$ dominate, so the comparison is between $2\sigma^{2}\|A_{2}\|_{\text{F}}^{2}$ for JME and $d^{2}\sigma^{4}\|A_{2}\|_{\text{F}}^{2}$ for PP. Consequently, at least for $\sigma\geq\sqrt{2}$, JME achieves privacy with less added noise than PP.
In the regime of low privacy ($\sigma\to 0$) in high dimension $(d\gg 1)$, PP can be expected to result in higher accuracy estimates than JME, because the terms involving $\sigma^{4}$ make only minor contributions.
For settings with general noise shaping matrix, we cannot provide an exact comparison between PP and JME, because the $\sup_{X}$-term in Equation 8 has no closed-form solution. Instead, we derive upper and lower bounds (Lemma C.1 in the Appendix), and provide a numeric comparison in Figure 2.
unbiased? symmetric noise? privacy-for-free?
Method 1st / 2nd moment 1st / 2nd moment 2nd moment recommended regime of use
JME (proposed) ✓/ ✓ ✓/ ✓ high privacy or low dimensionality
IME ✓/ ✓ ✓/ ✓ never (Pareto-inferior to JME)
CS ✓/ ✓ ✓/ ✓ never (Pareto-inferior to JME)
PP ✓/✗ ✓/ ✗ low privacy in high dimensions
PP (Debiased) ✓/✓ ✓/ ✗ low privacy in high dimensions
Table 2: Overview of the properties for the methods discussed in Section 3. Properties unbiased and symmetric noise apply to each entry, stated separately for first and second moment. privacy-for-free applies only for the second moment.
@@ 4 Applications
4.1 Private Gaussian density estimation
@@
The expected approximation error of $\widehat{\mu}$ is identical to the one in Theorem 3.3 with $A=V$ and $C_{1}=C_{2}=\text{I}$. The following Theorem establishes the approximation quality of the covariance estimate.
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}\|_{\text{F}}^{2}$ $=(c_{d}d^{2}+2d+2)\sigma^{2}H_{n,1}$ (11)
$\quad+d(d+1)\sigma^{4}H_{n,2},$
@@
Assume the same setting as for Theorem 4.1. Let $\widehat{\Sigma}_{\text{PP}}$ be the result of the above construction, where privacy is obtained by running PP with noise strength $\sigma$ and debiasing. Then, for the expected error of the covariance matrix estimate it holds:
$d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}$
$\qquad+2(d+1)\sigma^{2}H_{n,1}-2(d+1)\sigma^{2}H_{n,3},$
$\leq\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ (12)
$\leq d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}+2(d+1)\sigma^{2}H_{n,1}.$
@@
In the high privacy regime ($\sigma\gg 1$), the leading term for JME is $d(d+1)H_{n,2}\sigma^{4}$, and for PP it is $d(d+1)H_{n,1}\sigma^{4}$. Given that $H_{n,1}=O(\log n)$, while $H_{n,m}<\pi^{2}/6$ for $m\geq 2$, the error introduced by PP is logarithmically worse than JME. Figure 3 shows a numerical plot that confirms this observation. Note that our results match the lower bounds established by Kamath et al. (2020), who proved that private covariance estimation in the Frobenius norm requires $\Omega(d^{2})$ samples.
4.2 Private Adam optimization
The Adam optimizer (Diederik & Ba, 2015), has become a de-facto standard for optimization in deep learning. The defining property of Adam is its update rule, $\theta_{i}\leftarrow\theta_{i-1}-\alpha m_{i}/(\sqrt{v_{i}}+\epsilon)$, where $\alpha$ is a learning rate, and $m_{i}$ and $v_{i}$ are exponentially running averages of computed model gradients and componentwise squared model gradients, respectively. In the context of our work, these quantities correspond to a weighted first-moment vector and the diagonal of the weighted second-moment matrix.
Previous attempts to make Adam differentially private relied on postprocessing (Anil et al., 2022; Li et al., 2022), potentially with debiasing (Tang et al., 2024), i.e. they privatized the model gradients and derived the squared values from these. We demonstrate that JME’s approach of privatizing both quantities separately can be a competitive alternative. Algorithm 2 in the Appendix shows the pseudocode.
@@ 4.2 Private Adam optimization
$\sup_{X\in\mathcal{X}}\mathbb{E}_{Z}\|\widehat{D}_{\text{PP}}-D\|^{2}$ $=(2d\sigma^{4}+4\sigma^{2})\cdot\|A_{2}\|_{\text{F}}^{2}$
$\quad+\underbrace{d\sigma^{4}\cdot\big{\|}\sum_{i}(A_{2})_{i}\big{\|}_{2}^{2}}_{\text{bias}},$ (14)
where $A_{2}$ is the workload matrix obtained from the coefficients of Adam’s exponentially weighted averaging operations. The term marked ”bias” disappears if PP is debiased.
The proof of the first part follows directly from Theorem 3.3. The error for PP is computed separately in Lemma C.4 in the appendix.
This lemma shows that as long as $\sigma\geq 1$, the error introduced by JME is strictly lower than that of PP (i.e. classic DP-Adam (Li et al., 2022)) and even the debiased version of PP (i.e. DP-AdamBC (Tang et al., 2024)). Figure 4 illustrates the relation in a prototypical case.
(a) trivial factorization
(b) square root factorization
@@ 5 Experiments
Private Gaussian density estimation.
From a given data distribution, $p(x)=\mathcal{N}(\mu,\Sigma)$, we sample $n=200$ data points and use either JME or PP to form a private estimates, $\widehat{p}_{t}(x)=\mathcal{N}(\widehat{\mu}_{t},\widehat{\Sigma}_{t})$, at each step $t=1,\dots,n$, of the continuous release process. To ensure positive definiteness, we symmetrize JME’s estimated covariance matrices and project them onto the positive definite cone. As a postprocessing operation, this does not affect their privacy.
@@
Figure 5: Approximation quality of JME and PP for private Gaussian density estimation ($d=5,n=100$) as average and standard deviation over 1000 runs. In the high-privacy regime ($\sigma=2$), JME achieves lower KL divergence than PP after $\approx 10$ samples.
@@
Private model training with Adam.
We train a convolutional network on the CIFAR-10 dataset with DP-Adam, which is privatized either with JME or PP. Because gradients could have arbitrarily large number, we apply gradient clipping to a model-selected threshold in both methods. In addition, we include a heuristic baseline that uses the concatenate-and-split techniques in which not the norm of the gradient vector but the norm of the concatenation vector is clipped. While inferior to JME in the worst case, this might be beneficial for real data, so we include it in the experimental evalution.
@@ 6 Related Works
The problem of differentially private moments (and the related problem of covariance estimation) has a rich history of development, with optimal results known in the central model of privacy (Smith, 2011) as well as the local model of privacy (Duchi et al., 2013), and also in the worst-case setting. Sheffet (2019) proposes three differentially private algorithms for approximating the second-moment matrix, each ensuring positive definiteness. The related setting of covariance estimation has been studied in the worst-case setting when the data comes from a bounded $\ell_{2}$ ball by several works for approximate differential privacy (Achlioptas & McSherry, 2007; Blum et al., 2005; Dwork et al., 2014; Mangoubi & Vishnoi, 2022, 2023; Upadhyay, 2018) with the works of Amin et al. (2019) and Kapralov & Talwar (2013) presenting a pure differentially private algorithm.
Covariance estimation is also used as a subroutine in mean-estimation work under various distributional assumptions (Kamath, 2024). However, none of these approaches are directly applicable to the continual release model and they offer improvements over the Gaussian mechanism only in a very high privacy regime.

### v2

@@ 3.1 Comparison with alternative techniques
Besides JME, alternatives methods are possible that could be used to privately estimate the first and second moments. In this section, we introduce some of them and discuss their relation to JME.
Independent Moment Estimation (IME).
A straight-forward way to privately estimate first and second moments is to estimate and privatize both of them separately, where the necessary amount of noise is determined by the composition theorem of the Gaussian mechanism (abadi2016deep, ) (see Algorithm 4 in the appendix).
@@
Figure 1: Approximation errors for the 1st and 2nd moments across three methods (see text for details): $\lambda$-JME with $\lambda$, IME with varying $\alpha$ parameters, and CS with varying $\tau$ parameter. $\lambda$-JME Pareto-dominates the other two methods.
When applicable, CS is easy to implement and produces unbiased moment estimates. However, like IME, it has an unfavorable privacy-accuracy trade-off curve compared to JME, as formalized in the following theorem and shown in the appendix.
@@
Figure 1 visualizes these cases (Theorems 3.6 and 3.7) graphically in an exemplary setting $(d=10,n=100,C_{1}=C_{2}=I,A=E_{1},\zeta=1,\sigma_{\epsilon,\delta}=1/2)$.
Post-processing (PP).
Post-processing (PP) is another easy-to-use method for joint moment estimation. It has appeared in the literature dp_covariance_wishart_dist , at least in its naive form without the matrix factorization mechanism. For any $x_{i}$, it first computes a private estimate $\widehat{x_{t}}$ by adding sufficient noise to it. It then sets $\widehat{x_{t}\otimes x_{t}}:=\widehat{x_{t}}\otimes\widehat{x_{t}}$, which is automatically private by the postprocessing property of DP, and uses it to estimate the second moment matrix without additional privacy protection. Algorithm 6 in the appendix provides pseudocode.
(a) $d=1$, trivial
(b) $d=1$, square root
(c) $d=10^{3}$, trivial
(d) $d=10^{3}$, square root
@@
For given privacy parameters, PP and JME use the same noise strength toprivatize the first moment and therefore their estimates are of equal quality. For the second moment, their characteristics differ between the low privacy (small noise variance) and the high privacy (large noise variance) regime. To allow for a quantitative comparison, we derive characterizations of the approximation quality of PP.
@@
Assume the same setting as for Theorem 3.3, except that we compute the estimates $\widehat{Y}$ and $\widehat{S}_{\text{PP}}$ with the PP method. Let $Q=C_{1}^{-1}C_{1}^{-\top}$ and $E_{Q}=\operatorname{\operatorname{diag}}(Q)\operatorname{\operatorname{diag}}(Q)^{\top}$. Then, the expected approximation error of the second moment satisfies:
$\sup_{X\in\mathcal{X}}\mathbb{E}\|S-\widehat{S}_{\text{PP}}\|^{2}_{\text{F}}=16d(d+1)\sigma_{\epsilon,\delta}^{4}\zeta^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))$ (9)
$\quad\quad+8(d+1)\sigma_{\epsilon,\delta}^{2}\zeta^{2}\|C_{1}\|_{1\to 2}^{2}\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{\top}A_{2}\circ Q)XX^{T})\overbrace{+16d\sigma_{\epsilon,\delta}^{4}\zeta^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})}^{\text{bias}}.$
For debiased PP, the term marked "bias" does not occur.
@@
Assume that JME and PP use trivial factorizations, i.e., $C_{1}=C_{2}=I$. Then, JME’s expected approximation error for the second moment is
$\mathbb{E}\|S-\widehat{S}\|^{2}_{\text{F}}$ $=4c_{d}d^{2}\sigma_{\epsilon,\delta}^{2}\zeta^{4}\|A_{2}\|_{\text{F}}^{2}$ (10)
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}\|S-\widehat{S}_{\text{PP}}\|^{2}_{\text{F}}$ $=\big{(}16d(d+1)\zeta^{4}\sigma_{\epsilon,\delta}^{4}+8(d+1)\sigma_{\epsilon,\delta}^{2}\zeta^{2}\big{)}\|A_{2}\|_{\text{F}}^{2}\overbrace{+16d\sigma_{\epsilon,\delta}^{4}\zeta^{4}\|\sum_{i}(A_{2})_{[i,\cdot]}\|_{2}^{2}}^{\text{bias}},$
where $\sum_{i}(A_{2})_{[i,\cdot]}$ is the row-wise summation of $A_{2}$. For debiased PP, the term marked "bias" does not occur.
Proof. The proofs follow directly from Theorem 3.3 and Lemma 3.8 by observing that $\|C_{1}\|_{1\to 2}=\|C_{2}\|_{1\to 2}=1$. In dimension $d=1$, $\zeta=1$, JME has an error of $4c_{1}\sigma_{\epsilon,\delta}^{2}\|A_{2}\|_{\text{F}}^{2}$ versus $(32\sigma_{\epsilon,\delta}^{4}+16\sigma_{\epsilon,\delta}^{2})\|A_{2}\|_{\text{F}}^{2}$ for (debiased) PP. Since $c_{1}<1$, for one-dimensional data, JME’s error is always lower than PP’s (see Figure 2 for visual confirmation). In high dimensions ($d\gg 1$), the terms quadratic in $d$ dominate, so the comparison is between $8\sigma_{\epsilon,\delta}^{2}\|A_{2}\|_{\text{F}}^{2}$ for JME and $16d^{2}\sigma_{\epsilon,\delta}^{4}\|A_{2}\|_{\text{F}}^{2}$ for PP. Consequently, at least for $\sigma_{\epsilon,\delta}\geq\frac{1}{\sqrt{2}}$, JME achieves privacy with less added noise than PP.
In the regime of low privacy ($\sigma_{\epsilon,\delta}\to 0$) in high dimension $(d\gg 1)$, PP can be expected to result in higher accuracy estimates than JME, because the terms involving $\sigma_{\epsilon,\delta}^{4}$ make only minor contributions.
For settings with general noise shaping matrix, we cannot provide an exact comparison between PP and JME, because the $\sup_{X}$-term in Equation 9 has no closed-form solution. Instead, we derive upper and lower bounds (Lemma D.2), and provide a numeric comparison in Figure 2.
@@ 4 Applications
Private Gaussian density estimation
@@
The expected approximation error of $\widehat{\mu}$ is identical to the one in Theorem 3.3 with $A=V$ and $C_{1}=C_{2}=\text{I}$. The following Theorem establishes the approximation quality of the covariance estimate. In this and the following sections we denote $\sigma=2\sigma_{\epsilon,\delta}$, referring to the strength of the noise we add to the first moment for $\zeta=1$ and trivial factorization.
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}\|_{\text{F}}^{2}$ $=(c_{d}d^{2}+2d+2)\sigma^{2}H_{n,1}+d(d+1)\sigma^{4}H_{n,2},$ (11)
@@
Assume the same setting as for Theorem 4.1. Let $\widehat{\Sigma}_{\text{PP}}$ be the result of the above construction, where privacy is obtained by running PP with noise strength $\sigma$ and debiasing. Let $S(n,d,\sigma):=d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}+2(d+1)\sigma^{2}H_{n,1}$. Then, for the expected error of the covariance matrix estimate it holds:
$S(n,d,\sigma)-2(d+1)\sigma^{2}H_{n,3}\leq\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}\leq S(n,d,\sigma).$ (12)
@@
In the high privacy regime ($\sigma\gg 1$), the leading term for JME is $d(d+1)H_{n,2}\sigma^{4}$, and for PP it is $d(d+1)H_{n,1}\sigma^{4}$. Given that $H_{n,1}=O(\log n)$, while $H_{n,m}<\pi^{2}/6$ for $m\geq 2$, the error introduced by PP is logarithmically worse than JME. Figure 6 shows a numerical plot that confirms this observation. Note that our results match the lower bounds established by G. Kamath 2020 (dp_covariance_lower_bound, ), who proved that private covariance estimation in the Frobenius norm requires $\Omega(d^{2})$ samples.
Private Adam optimization
The Adam optimizer (kingma2014adam, ), has become a de-facto standard for optimization in deep learning. The defining property of Adam is its update rule, $\theta_{i}\leftarrow\theta_{i-1}-\alpha m_{i}/(\sqrt{v_{i}}+\epsilon)$, where $\alpha$ is a learning rate, and $m_{i}$ and $v_{i}$ are exponentially running averages of computed model gradients and componentwise squared model gradients, respectively. In the context of our work, these quantities correspond to a weighted first-moment vector and the diagonal of the weighted second-moment matrix.
Previous attempts to make Adam differentially private relied on postprocessing (anil2021large, ; li2022large, ), potentially with debiasing (dp_adam_is_not_adam, ), i.e. they privatized the model gradients and derived the squared values from these. We demonstrate that JME’s approach of privatizing both quantities separately can be a competitive alternative. Algorithm 2 in the Appendix shows the pseudocode.
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}_{Z}\|\widehat{D}_{\text{PP}}-D\|^{2}$ $=(2d\sigma^{4}+4\sigma^{2})\cdot\|A_{2}\|_{\text{F}}^{2}\ \overbrace{+d\sigma^{4}\cdot\big{\|}\sum_{i}(A_{2})_{i}\big{\|}_{2}^{2}}^{\text{bias}},$ (14)
where $A_{2}$ is the workload matrix obtained from the coefficients of Adam’s exponentially weighted averaging operations. The term marked "bias" disappears if PP is debiased.
The proof of the first part follows directly from Theorem 3.3. The error for PP is computed separately in Lemma D.5 in the appendix.
This lemma shows that as long as $\sigma\geq 1$, the error introduced by JME is strictly lower than that of PP (i.e. classic DP-Adam (li2022large, )) and even the debiased version of PP (i.e. DP-AdamBC (dp_adam_is_not_adam, )). Figure 5 illustrates the relation in a prototypical case.
@@ 5 Experiments
Private Gaussian density estimation. From a given data distribution, $p(x)=\mathcal{N}(\mu,\Sigma)$, we sample $n=200$ data points and use either JME or PP to form a private estimates, $\widehat{p}_{t}(x)=\mathcal{N}(\widehat{\mu}_{t},\widehat{\Sigma}_{t})$, at each step $t=1,\dots,n$, of the continuous release process. To ensure positive definiteness, we symmetrize JME’s estimated covariance matrices and project them onto the positive definite cone. As a postprocessing operation, this does not affect their privacy.
@@
(a) $d=5,n=100,\sigma=2$
(b) $d=10,n=200,\sigma=4$
Figure 3: Approximation quality of JME and PP for private Gaussian density estimation as average and standard deviation over 1000 runs. In the high-privacy regime ($\sigma=2$ and $\sigma=4$), JME achieves lower KL divergence than PP after $\approx 10$ samples.
@@
Private model training with Adam. We train a convolutional network on the CIFAR-10 dataset with DP-Adam, which is privatized either with JME or PP. Because gradients could have arbitrarily large norm, we apply gradient clipping to a model-selected threshold in both methods. In addition, we include a heuristic baseline that uses the concatenate-and-split techniques in which not the norm of the gradient vector but the norm of the concatenation vector is clipped. While inferior to JME in the worst case, this might be beneficial for real data, so we include it in the experimental evalution.
@@ 6 Related works
The problem of differentially private moments (and the related problem of covariance estimation) has a rich history of development, with optimal results known in the central model of privacy (smith2011privacy, ) as well as the local model of privacy (duchi2013local, ), and also in the worst-case setting. Sheffet (dp_covariance_wishart_dist, ) proposes three differentially private algorithms for approximating the second-moment matrix, each ensuring positive definiteness. The related setting of covariance estimation has been studied in the worst-case setting when the data comes from a bounded $\ell_{2}$ ball by several works for approximate differential privacy (achlioptas2007fast, ; blum2005practical, ; dwork2014analyze, ; mangoubi2022re, ; mangoubi2023private, ; upadhyay2018price, ) with the works of Amin et al. (dp_covariance, ) and Kapralov & Talwar (kapralov2013differentially, ) presenting a pure differentially private algorithm. Covariance estimation is also used as a subroutine in mean-estimation work under various distributional assumptions (kamath2024broader, ). However, none of these approaches are directly applicable to the continual release model and they offer improvements over the Gaussian mechanism only in a very high privacy regime. Precise sensitivity analysis, as a technique to improve differential privacy guarantees, has also been intensively used in the literature (lebeda2024correlated, ; wilkins2024exact, ; joseph2024privately, ; lebeda2025better, ). The matrix mechanism has also gained a lot of attention recently due to its application of continually releasing prefix sums in private online optimization BandedMatrix ; Kairouz ; efficient_streaming ; henzinger2025improved ; fichtenberger2023constant ; henzinger2023almost ; kalinin2025back ; henzinger2025binned .